OpenStudy (anonymous):

Matt has eight times as many stickers as David. how many stickers myst matt give david so that they will each have 90 stickers?

OpenStudy (mertsj):

m+d=180 m=8d

OpenStudy (mertsj):

Solve that system. Then you will know how many stickers each has. Then you can figure out how many to subtract from Matt's and add to David's so that they each have 90.

OpenStudy (anonymous):

I am confused. Dont knolw how to solve the system.

OpenStudy (mertsj):

replace m in the first equation with 8d from the second equation and solve that equation.

OpenStudy (anonymous):

Sorry, I am lost and dont understand this probelm...

OpenStudy (mertsj):

What is the first equation I wrote?

OpenStudy (anonymous):

m+d=180

OpenStudy (mertsj):

Do you see that the second equation says that m = 8d?

OpenStudy (anonymous):

yes....sorry I am completely dumb with math

OpenStudy (mertsj):

so take the equation m+d = 180 and erase the m and put 8d there because m = 8d. Now what do you have?

OpenStudy (anonymous):

8d=180

OpenStudy (anonymous):

would it be 44 stickers

OpenStudy (anonymous):

or 23

OpenStudy (mertsj):